A right circular cylindrical vessel is exactly one-quarter full. If 7 liters of liquid are added, the vessel will be exactly three-fifths full. What is the total capacity of the vessel, in liters?
Answer by mananth(16946)

let the total capacity of vessel be x
Now it is x/4 full
7 liters are added
x/4 +7
x/4 +7 = 3x/5
x/4-3x/5 = -7
LCD = 20
Multiply equation by LCD
5x-12x=20*-7
-7x=-140
/-7
x=20
the full capacity is 20 liters

m.ananth@hotmail.ca